### v1

@@ 5.6 Comparisons with Advanced Metrics
Backbone Reward B-4 M C S RefCLIP-S RefPAC-S++ CLIP-S PAC-S++ Rep-1 Rep-2 Rep-3 Rep-4 %Incorrect
- 33.1 28.2 112.4 20.5 0.804 0.794 0.755 0.712 1.468 0.091 0.017 0.005 0.3
CIDEr 40.4 29.4 129.6 21.6 0.806 0.799 0.751 0.714 1.318 0.038 0.006 0.004 24.7
CLIP-S 12.1 23.5 1.1 20.0 0.767 0.776 0.844 0.744 12.226 4.736 1.884 0.795 99.2
PAC-S++ 19.4 27.1 36.3 22.4 0.801 0.795 0.813 0.755 5.129 1.431 0.544 0.229 0.7
RefCLIP-S 26.3 27.6 92.5 21.4 0.829 0.807 0.799 0.735 2.571 0.626 0.236 0.103 0.3
ViT-B/32 RefPAC-S++ 30.5 28.5 109.1 22.2 0.822 0.811 0.784 0.740 1.791 0.247 0.069 0.026 0.3
- 34.8 29.9 119.4 22.5 0.802 0.708 0.749 0.708 1.469 0.064 0.008 0.002 0.3
CIDEr 43.6 30.8 143.3 23.2 0.809 0.804 0.750 0.713 1.432 0.047 0.005 0.002 32.3
CLIP-S 13.1 24.6 1.4 20.0 0.782 0.780 0.840 0.736 11.225 4.447 2.08 1.012 34.8
PAC-S++ 20.9 28.0 51.8 23.9 0.806 0.797 0.812 0.751 4.157 0.974 0.33 0.129 1.3
RefCLIP-S 27.8 28.8 101.9 23.3 0.833 0.811 0.800 0.734 2.161 0.386 0.13 0.046 0.7
ViT-L/14 RefPAC-S++ 32.5 29.6 118.9 23.5 0.826 0.814 0.782 0.736 1.468 0.145 0.037 0.011 0.9
@@
5.7 PAC-Score++ for RL-based Captioning Fine-tuning
We then evaluate the effectiveness of the proposed PAC-S++ metric when employed as reward for fine-tuning a captioning model, using the fine-tuning strategy described in Sec. 4. In this setting, we compare our metric in both its reference-free and reference-based version respectively against CLIP-S and RefCLIP-S. For completeness, we also report the results of the model trained with cross-entropy loss only (i.e. without reinforcement learning) and using the standard CIDEr score as reward. To evaluate generated captions, we employ a combination of traditional metrics, like BLEU [43], METEOR [4], CIDEr [63], and SPICE [3], and more recent ones such as CLIP-S and the proposed PAC-S++ metric, considering in both cases reference-based and reference-free settings. Additionally, we introduce novel metrics to assess the grammatical correctness of the generated captions, which is crucial especially when directly optimizing CLIP-based scores. Specifically, we measure the average number of repeated $n$-grams (Rep-$n$) and the percentage of captions ending with undesirable words like prepositions, conjunctions, or determiners (%Incorrect).
In-domain Evaluation. Captioning results on the COCO test set are reported in Table 7. Notably, although CLIP remains an excellent model for aligning bag-of-words with visual input, it disregards syntax and logical connections among words within captions. On the contrary, despite sharing the same architecture, our proposal mitigates this issue, favouring the use of PAC-S++ as a reward metric in a captioning model. In particular, directly optimizing CLIP-S leads to protracted and repetitive captions, as demonstrated by the lower scores in terms of standard reference-based metrics and grammar measures. In contrast, PAC-S++ significantly stabilizes the fine-tuning process, yielding significant enhancements in reference-based metrics (e.g.  36.3 and 51.8 CIDEr points using PAC-S++ with ViT-B/32 and ViT-L/14 features vs. 1.1 and 1.4 obtained by CLIP-S). Concurrently, it enables the generation of semantically rich and grammatically correct captions that better correlate with human-generated content. This phenomenon is particularly notable in repetitiveness metrics, where the average number of repeated 1-grams in the generated captions decreases from $11.225$ to $4.157$, when using VIT-L/14 as visual backbone.
@@ 5.7 PAC-Score++ for RL-based Captioning Fine-tuning
Table 8: Captioning results in terms of reference-based and reference-free evaluation metrics on nocaps and VizWiz validation sets.
@@
Backbone Reward C CLIP-S PAC-S++ RefCLIP-S RefPAC-S++ C CLIP-S PAC-S++ RefCLIPS RefPAC-S++
- 67.6 0.686 0.694 0.699 0.733 27.8 0.655 0.675 0.691 0.729
CIDEr 76.2 0.695 0.703 0.709 0.741 28.9 0.663 0.686 0.704 0.739
CLIP-S 1.6 0.780 0.726 0.675 0.724 1.1 0.735 0.703 0.686 0.722
PAC-S++ 34.6 0.751 0.743 0.713 0.748 17.5 0.721 0.729 0.717 0.751
Ref-CLIP-S 64.0 0.736 0.724 0.734 0.753 25.0 0.703 0.708 0.723 0.747
ViT-B/32 RefPAC-S++ 73.1 0.724 0.729 0.728 0.758 29.4 0.694 0.715 0.723 0.758
- 75.2 0.698 0.704 0.710 0.743 35.0 0.655 0.679 0.701 0.740
CIDEr 91.3 0.698 0.711 0.718 0.755 39.6 0.667 0.683 0.722 0.751
CLIP-S 2.1 0.791 0.741 0.705 0.746 1.6 0.727 0.703 0.711 0.741
PAC-S++ 49.1 0.769 0.754 0.735 0.764 26.1 0.713 0.723 0.726 0.759
Ref-CLIP-S 79.0 0.756 0.742 0.756 0.774 35.0 0.705 0.708 0.738 0.761
ViT-L/14 RefPAC-S++ 89.8 0.741 0.744 0.750 0.776 41.3 0.695 0.715 0.737 0.770
Out-of-domain Evaluation. Finally, we evaluate the out-of-domain performance of our model on the nocaps [2] and VizWiz [15] datasets, both of which present distinct image descriptions compared to the COCO dataset used for training. Specifically, the nocaps dataset, which is designed for the novel object captioning task, includes image-caption pairs featuring objects not present in the COCO training set. In contrast, VizWiz consists of images taken by visually impaired individuals, often showcasing challenging perspectives, such as close-up shots or unconventional viewpoints. The results, summarized in Table 8, are evaluated using both reference-free and reference-based metrics.
Also in these challenging settings, our approach demonstrates greater semantic richness while preserving fluidity and grammatical correctness in text generation. This behaviour is not observed when CLIP-S is used as reward. Specifically, although the use of CLIP-S results in high scores on learnable metrics, the values of traditional metrics remain notably low. For instance, on the nocaps dataset and using VIT-L/14 as visual backbone, the CIDEr score drops dramatically from 49.1 points when using PAC-S++ as reward to just 2.1 points with CLIP-S as reward, further highlighting the advantages of our proposed metric for training captioning models.

### v2

@@ B-4 M C S RefCLIP-S Polos RefPAC-S++ CLIP-S PAC-S++
5.6 System-Level Correlation
In Table 6, we assess whether well established captioning metrics and our proposed one effectively evaluate captions generated by popular existing captioning model and modern LLM-based architectures. Among the standard captioning models, we include Show and Tell [72] and Show, Attend and Tell [76], Up-Down [4], AoANet [23], $\mathcal{M}^{2}$ Transformer [11], X-Transformer [50], DLCT [46], and the more recently proposed COS-Net model [40] that incorporates CLIP features. In addition to reporting evaluations on traditional captioning models, we also include recent LLM-based captioning models, like BLIP-2 [37] [^2] Note that for BLIP-2 we employ the model fine-tuned on COCO image-caption pairs., IDEFICS-9B [32], LLaVA-1.5-7B [43] and ShareGPTV4 [8]. We evaluate PAC-S++ scores against both standard metrics, such as BLEU-4, METEOR, CIDEr, and SPICE, and more recent ones like CLIP-S and Polos. To assess performance, captioning models are compared to a human baseline, where, for each sample, one human-annotated sentence (randomly selected from the five available in the COCO dataset) is used as the candidate caption.
As shown, PAC-S++ effectively evaluates human-written captions, assigning scores comparable to those of recent state-of-the-art models like COS-Net and certain LLM-based approaches. In contrast, standard metrics such as METEOR and CIDEr often rank human captions below those generated by weaker models like Show and Tell or Up-Down. This limitation is particularly evident when evaluating captions from LLM-based models – especially longer ones, such as those produced by ShareGPT4V – where traditional metrics like CIDEr tend to struggle, as shown in the qualitative examples reported in Fig. 4. In these scenarios, reference-free metrics such as CLIP-S and PAC-S++ show a stronger ability to handle longer, LLM-generated captions and provide more reliable evaluations.
5.7 Ablation Studies
Effect of the Scaling Factor $w$. The scaling factor, denoted by $w$ in Eq. 4, is used to adjust the scale of the final metric. This adjustment is made to enhance the numerical readability without impacting the ranking of the results. Notably, CLIP-S proposes this analysis and sets $w=2.5$. In our case, due to the different score distributions, we use different $w$ values when using different backbones, as shown in Fig. 5. Specifically, all experiments featuring the ViT-B/32 backbone employ $w$ set at 2.52.52.5, while for ViT-L/14, the value of $w$ is set to 333.
Table 7: Ablation study results of PAC-S++/ RefPAC-S++, using different hyperparameters and synthetic data generators. For each experiment, we report in the last column the averaged improvement compared to the previous version of our metric (i.e. PAC-S [61]), reported in the first row.
Synthetic Data Flickr8k-Exp Flickr8k-CF Composite VATEX-EVAL Pascal-50S FOIL ActivityNet-FOIL
LoRA $r$ $\lambda_{v}$ $\lambda_{t}$ Visual Textual Kendall $\tau_{c}$ Kendall $\tau_{b}$ Kendall $\tau_{c}$ Kendall $\tau_{b}$ Accuracy Accuracy Accuracy $\bar{\Delta}$
- 0.05 0.1 SDv1.5 BLIP 53.9 / 55.5 36.0 / 37.6 55.7 / 57.3 25.1 / 31.4 82.4 / 84.7 89.9 / 93.7 90.1 / 93.5
Effect of Varying LoRA $r$
2 0.1 0.001 SDv1.5 BLIP 54.3 / 55.6 36.9 / 38.0 58.2 / 59.1 27.4 / 32.3 82.2 / 84.4 90.1 / 93.5 91.1 / 93.6 +0.71
4 0.1 0.001 SDv1.5 BLIP 54.5 / 55.7 37.0 / 37.9 58.3 / 59.1 28.1 / 32.2 82.3 / 84.5 90.2 / 93.5 91.0 / 93.4 +0.78
8 0.1 0.001 SDv1.5 BLIP 54.5 / 55.6 36.9 / 38.0 58.3 / 59.2 27.2 / 32.1 82.0 / 84.5 90.0 / 93.6 90.7 / 93.5 +0.66
16 0.1 0.001 SDv1.5 BLIP 54.3 / 55.5 37.0 / 37.9 58.5 / 59.3 27.8 / 32.3 81.8 / 84.5 90.1 / 93.6 90.9 / 93.7 +0.74
Effect of Varying $\lambda_{v}$ and $\lambda_{t}$
4 0.001 0.001 SDv1.5 BLIP 54.6 / 55.8 37.1 / 38.0 57.4 / 58.2 27.9 / 32.4 81.9 / 84.5 90.1 / 93.6 90.4 / 93.3 +0.60
4 0.05 0.001 SDv1.5 BLIP 54.7 / 55.8 36.9 / 37.9 58.3 / 59.1 27.5 / 32.0 82.1 / 84.5 90.3 / 93.7 91.0 / 93.7 +0.76
4 0.1 0.001 SDv1.5 BLIP 54.5 / 55.7 37.0 / 37.9 58.3 / 59.1 28.1 / 32.2 82.3 / 84.5 90.2 / 93.5 91.0 / 93.4 +0.78
4 0.5 0.001 SDv1.5 BLIP 54.4 / 55.8 36.8 / 37.8 57.9 / 58.9 27.6 / 31.9 81.8 / 84.1 89.8 / 93.6 90.4 / 93.4 +0.53
4 0.1 0.05 SDv1.5 BLIP 54.7 / 55.8 37.0 / 38.0 58.0 / 58.9 27.7 / 32.2 81.9 / 84.4 90.4 / 93.6 90.9 / 93.5 +0.73
4 0.1 0.1 SDv1.5 BLIP 54.6 / 55.8 37.0 / 38.0 57.5 / 58.5 27.7 / 32.0 82.3 / 84.3 89.9 / 93.4 91.3 / 93.3 +0.63
Synthetic Data Contribution
4 0.1 0.001 - - 53.9 / 54.9 36.7 / 37.7 57.5 / 58.4 26.7 / 31.9 81.7 / 83.9 89.8 / 93.2 90.1 / 93.2 +0.20
4 0.1 0.001 - BLIP 54.4 / 55.4 36.8 / 37.8 57.3 / 58.3 26.7 / 32.0 82.1 / 84.4 89.8 / 93.3 90.9 / 93.6 +0.43
4 0.1 0.001 SDv1.5 - 54.2 / 55.2 37.0 / 37.9 57.9 / 58.7 27.0 / 32.1 82.0 / 84.4 89.9 / 93.3 90.4 / 93.7 +0.49
4 0.1 0.001 SDv1.5 BLIP 54.5 / 55.7 37.0 / 37.9 58.3 / 59.1 28.1 / 32.2 82.3 / 84.5 90.2 / 93.5 91.0 / 93.4 +0.78
Effect of Varying Synthetic Data Sources
4 0.1 0.001 SDv1.5 BLIP 54.5 / 55.7 37.0 / 37.9 58.3 / 59.1 28.1 / 32.2 82.3 / 84.5 90.2 / 93.5 91.0 / 93.4 +0.78
4 0.1 0.001 SDv1.5 IDEFICS-3 54.2 / 55.4 36.6 / 37.8 58.0 / 58.9 26.9 / 31.9 81.9 / 84.2 89.6 / 93.3 90.9 / 93.8 +0.47
4 0.1 0.001 SDv1.5 LLaMA-3.2 54.6 / 55.7 36.9 / 38.0 57.9 / 58.7 27.1 / 31.9 82.1 / 84.3 90.2 / 93.7 90.9 / 93.8 +0.64
4 0.1 0.001 SDv3.5 BLIP 54.7 / 56.0 36.8 / 37.9 58.1 / 58.9 27.9 / 31.9 82.0 / 84.4 90.2 / 93.4 90.9 / 93.6 +0.71
4 0.1 0.001 FLUX.1 BLIP 54.8 / 56.0 37.0 / 38.0 58.2 / 59.0 28.0 / 32.0 81.8 / 84.4 90.2 / 93.4 91.1 / 93.6 +0.76
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Low-Rank Analysis. All the analyses conducted so far employ the PAC-S++ version with low-rank adaptation. In Table 7, we investigate the effect of different ranks (i.e. 2, 4, 8, 16) across the selected datasets. Overall, the best performance are achieved with a rank of 4, in both reference-free and reference-based settings.
Notably, comparing the results of the previous version of our metric (i.e. PAC-S [61], first row), in which only the last visual and textual projections of the model are fine-tuned, the version of our metric with LoRA consistently outperforms the original version regardless of the rank, with the only exception of Pascal-50S and FOIL datasets. For example, employing PAC-S++ with the CLIP ViT-B/32 backbone fine-tuned with LoRA yields superior results compared to its counterpart without LoRA, achieving 58.3 on the Composite dataset (+2.6) and 28.1 on the VATEX-EVAL dataset (+3.0). This demonstrates the effectiveness of this strategy, even in the context of video settings.
Choice of Hyperparameters. Subsequent to determining the optimal rank dimension as $r=4$, we also conduct a comprehensive grid search to determine the optimal values of the $\lambda_{v}$ and $\lambda_{t}$ hyperparameters for our model across multiple datasets. The results are summarized in Table 7. While the results show notable variation in performance depending on the dataset, we observe that the configuration corresponding $\lambda_{v}=0.1$ and $\lambda_{t}=0.001$ consistently yields strong improvements across most datasets. Accordingly, we adopt this setting as our final configuration, which is used in the loss function defined in Eq. 3.
Table 8: Human correlation and accuracy results when varying visual and textual backbones. The best results for each backbone are highlighted in bold.
Flickr8k-Exp Flickr8k-CF Composite Pascal-50S FOIL
Backbone Kendall $\tau_{b}$ Kendall $\tau_{c}$ Kendall $\tau_{b}$ Kendall $\tau_{c}$ Kendall $\tau_{b}$ Kendall $\tau_{c}$ Accuracy Accuracy
CLIP-S 52.6 53.0 35.2 18.2 51.3 55.4 81.7 90.9
PAC-S 55.1 55.5 36.8 19.0 52.3 56.5 82.2 91.9
PAC-S++ 57.0 57.4 38.5 19.9. 57.3 62.0 82.4 93.6
RefCLIP-S 54.1 54.5 36.5 18.9 51.9 56.1 84.3 94.0
RefPAC-S 56.7 57.1 37.7 19.5 53.1 57.2 85.0 95.3
CLIP ViT-L/14 RefPAC-S++ 57.5 57.9 38.8 20.1 60.0 61.6 85.4 95.1
CLIP-S 0.49 0.49 -1.1 -0.6 26.1 28.2 62.2 52.5
PAC-S 58.8 54.0 39.7 19.9 54.5 58.9 81.7 92.3
PAC-S++ 61.3 42.7 48.5 19.7 58.8 61.5 81.8 94.1
RefCLIP-S 33.3 33.6 18.8 9.7 30.9 33.4 73.0 69.8
RefPAC-S 59.0 54.2 40.0 20.0 55.1 59.5 83.1 93.6
SigLIP ViT-L/14 RefPAC-S++ 61.4 42.8 48.5 19.7 58.8 61.6 82.7 94.4
CLIP-S 4.6 4.5 1.2 0.6 19.6 21.2 59.3 61.0
PAC-S 4.5 4.1 1.1 0.5 8.6 8.6 54.9 41.2
PAC-S++ 60.1 47.4 42.6 19.3 59.1 62.6 81.0 93.4
RefCLIP-S 4.7 4.6 1.2 0.4 29.9 32.4 72.7 77.3
RefPAC-S 4.5 4.2 1.0 0.5 9.7 9.7 55.3 43.4
SigLIP2 ViT-L/14 RefPAC-S++ 60.2 47.5 42.6 19.3 59.2 62.7 81.6 93.9
Synthetic Data Generator Analysis. To assess the overall contribution of synthetic data and the impact of different text and image generators, we conduct an ablation study, always reported in Table 7. We start by analyzing the contribution of each generator. Specifically, we conduct experiments without synthetic data (i.e. only using pairs from COCO during fine-tuning), as well as analyze the effects of removing only the textual or the visual positive augmentations. The averaged improvements compared to the previous version of the metric highlight that both visual and textual synthetic data contribute to enhanced performance compared to the version without any positive augmentation. Furthermore, when both augmentations are applied simultaneously, further performance improvement is observed.
As additional analysis, we replace synthetic data generators with more recent models to assess their impact on performance. For text generation, we compare our default setup, which employs BLIP [38], with LLaMA 3.2 [17] and IDEFICS-3 [31] as caption generators. For image generation, instead, we assess Stable Diffusion v3.5 (SDv3.5) [^3] https://huggingface.co/stabilityai/stable-diffusion-3.5-large and FLUX [^4] https://huggingface.co/black-forest-labs/FLUX.1-dev as alternatives to our baseline, Stable Diffusion V1.5 (SDv1.5). Our analysis indicates that the original configuration remains the most effective overall. In particular, on the text side, substituting newer models does not yield noticeable performance improvements. This can be due to the behavior of modern text generators, which tend to produce longer and more verbose captions. CLIP textual encoder often truncates these, thereby limiting their effective contribution. On the image side, although newer generators may offer improved visual quality, results are comparable with the initial version using SDv1.5 images, overall confirming the effectiveness of our positive-augmented fine-tuning strategy.
Effect of Changing Backbone. We investigate the impact of changing the backbone architecture by replacing the standard CLIP visual and textual encoders with those from SigLIP [80] and SigLIP2 [69], which differ from CLIP in both training data and training strategies. To ensure a fair comparison, we use the same ViT-L/14 backbone for all three models and we report results in Table 8. For each, we evaluate both the reference-free and reference-based variants, applying the standard CLIP-S setup (analogous to a zero-shot setting), as well as PAC-S and our PAC-S++ training strategies. Our results show that CLIP-S with the CLIP ViT-L/14 backbone, while underperforming compared to PAC-S and PAC-S++, still achieves competitive results. However, when using SigLIP or SigLIP2, CLIP-S performance significantly degrades, particularly in terms of human correlation, pairwise ranking, and hallucination detection. This suggests that the zero-shot capabilities of SigLIP-like models are less aligned with the image captioning evaluation tasks. By applying the PAC-S training strategy to SigLIP, we observe substantial improvements, often surpassing even the performance of PAC-S++ with the CLIP ViT-L/14 backbone across several datasets. Further gains are achieved using the PAC-S++ strategy. For instance, with the SigLIP backbone, RefPAC-S++ achieves $\tau_{c}$ scores of 61.4 and 58.8 on the Flickr8k-Expert and Composite datasets, respectively, which are significantly higher than the CLIP-based version. These results highlight the importance of both the model architecture and the fine-tuning strategy in achieving robust performance across captioning evaluation benchmarks.
5.8 PAC-Score++ for RL-based Captioning Fine-tuning
[…]
In-domain Evaluation. Captioning results on the COCO test set are reported in Table 9. Notably, although CLIP remains an excellent model for aligning bag-of-words with visual input, it disregards syntax and logical connections among words within captions. On the contrary, despite sharing the same architecture, our proposal mitigates this issue, favouring the use of PAC-S++ as a reward metric in a captioning model. In particular, directly optimizing CLIP-S leads to protracted and repetitive captions, as demonstrated by the lower scores in terms of standard reference-based metrics and grammar measures. In contrast, PAC-S++ significantly stabilizes the fine-tuning process, yielding significant enhancements in reference-based metrics (e.g.  36.3 and 51.8 CIDEr points using PAC-S++ with ViT-B/32 and ViT-L/14 features vs. 1.1 and 1.4 obtained by CLIP-S). Concurrently, it enables the generation of semantically rich and grammatically correct captions that better correlate with human-generated content. This phenomenon is particularly notable in repetitiveness metrics, where the average number of repeated 1-grams in the generated captions decreases from 11.22511.22511.225 to 4.1574.1574.157, when using VIT-L/14 as visual backbone.
@@ 5.8 PAC-Score++ for RL-based Captioning Fine-tuning
Backbone Reward B-4 M C S RefCLIP-S Polos RefPAC-S++ CLIP-S PAC-S++ Rep-1 Rep-2 Rep-3 Rep-4 %Incorrect
- 33.1 28.2 112.4 20.5 0.804 0.651 0.794 0.755 0.712 1.468 0.091 0.017 0.005 0.3
CIDEr 40.4 29.4 129.6 21.6 0.806 0.651 0.799 0.751 0.714 1.318 0.038 0.006 0.004 24.7
CLIP-S 12.1 23.5 1.1 20.0 0.767 0.635 0.776 0.844 0.744 12.226 4.736 1.884 0.795 99.2
PAC-S++ 19.4 27.1 36.3 22.4 0.801 0.658 0.795 0.813 0.755 5.129 1.431 0.544 0.229 0.7
RefCLIP-S 26.3 27.6 92.5 21.4 0.829 0.679 0.807 0.799 0.735 2.571 0.626 0.236 0.103 0.3
ViT-B/32 RefPAC-S++ 30.5 28.5 109.1 22.2 0.822 0.677 0.811 0.784 0.740 1.791 0.247 0.069 0.026 0.3
- 34.8 29.9 119.4 22.5 0.802 0.078 0.708 0.749 0.708 1.469 0.064 0.008 0.002 0.3
CIDEr 43.6 30.8 143.3 23.2 0.809 0.668 0.804 0.750 0.713 1.432 0.047 0.005 0.002 32.3
CLIP-S 13.1 24.6 1.4 20.0 0.782 0.656 0.780 0.840 0.736 11.225 4.447 2.08 1.012 34.8
PAC-S++ 20.9 28.0 51.8 23.9 0.806 0.675 0.797 0.812 0.751 4.157 0.974 0.33 0.129 1.3
RefCLIP-S 27.8 28.8 101.9 23.3 0.833 0.700 0.811 0.800 0.734 2.161 0.386 0.13 0.046 0.7
ViT-L/14 RefPAC-S++ 32.5 29.6 118.9 23.5 0.826 0.702 0.814 0.782 0.736 1.468 0.145 0.037 0.011 0.9
@@
Out-of-domain Evaluation. Finally, we evaluate the out-of-domain performance of our model on the nocaps [2] and VizWiz [18] datasets, both of which present distinct image descriptions compared to the COCO dataset used for training. Specifically, the nocaps dataset, which is designed for the novel object captioning task, includes image-caption pairs featuring objects not present in the COCO training set. In contrast, VizWiz consists of images taken by visually impaired individuals, often showcasing challenging perspectives, such as close-up shots or unconventional viewpoints. The results, summarized in Table 10, are evaluated using both reference-free and reference-based metrics. Also in these challenging settings, our approach demonstrates greater semantic richness while preserving fluidity and grammatical correctness in text generation. This behavior is not observed when CLIP-S is used as reward. Specifically, although the use of CLIP-S results in high scores on learnable metrics, the values of traditional metrics remain notably low. For instance, on the nocaps dataset and using VIT-L/14 as visual backbone, the CIDEr score drops from 49.1 points when using PAC-S++ as reward to just 2.1 points with CLIP-S as reward, further highlighting the advantages of our proposed metric for training captioning models.
Table 10: Captioning results in terms of reference-based and reference-free metrics on nocaps and VizWiz validation sets.
@@
Backbone Reward C CLIP-S PAC-S++ RefCLIP-S Polos RefPAC-S++ C CLIP-S PAC-S++ RefCLIPS Polos RefPAC-S++
- 67.6 0.686 0.694 0.699 0.432 0.733 27.8 0.655 0.675 0.691 0.360 0.729
CIDEr 76.2 0.695 0.703 0.709 0.418 0.741 28.9 0.663 0.686 0.704 0.345 0.739
CLIP-S 1.6 0.780 0.726 0.675 0.467 0.724 1.1 0.735 0.703 0.686 0.403 0.722
PAC-S++ 34.6 0.751 0.743 0.713 0.485 0.748 17.5 0.721 0.729 0.717 0.434 0.751
Ref-CLIP-S 64.0 0.736 0.724 0.734 0.475 0.753 25.0 0.703 0.708 0.723 0.405 0.747
ViT-B/32 RefPAC-S++ 73.1 0.724 0.729 0.728 0.474 0.758 29.4 0.694 0.715 0.723 0.412 0.758
- 75.2 0.698 0.704 0.710 0.473 0.743 35.0 0.655 0.679 0.701 0.407 0.740
CIDEr 91.3 0.698 0.711 0.718 0.438 0.755 39.6 0.667 0.683 0.722 0.365 0.751
CLIP-S 2.1 0.791 0.741 0.705 0.508 0.746 1.6 0.727 0.703 0.711 0.427 0.741
PAC-S++ 49.1 0.769 0.754 0.735 0.536 0.764 26.1 0.713 0.723 0.726 0.474 0.759
Ref-CLIP-S 79.0 0.756 0.742 0.756 0.528 0.774 35.0 0.705 0.708 0.738 0.456 0.761
ViT-L/14 RefPAC-S++ 89.8 0.741 0.744 0.750 0.530 0.776 41.3 0.695 0.715 0.737 0.468 0.770